How do I write the equation for the question below, and then how do I solve it.
Questions: One angle is 3 times another. If the two angles are complimentary, what are the angle measures?

Answer by stanbon(75887)

One angle is 3 times another. If the two angles are complimentary, what are the angle measures?
------------
x = 3y
x+y = 90
-------------
solve by substitution:
3y+y = 90
4y = 90
y = 22.5 degrees
----------------
Substitute into x = 3y to get
x = 3*22.5 = 67.5 degrees
==============================
Cheers,
Stan H.